Farmer Fred has goats and chickens in his fenced barnyard. Tiny Tim can only see under the fence and he counts 72 feet. Tall Tony can see over the fence and he counts 26 heads. How many chickens does Farmer Fred have in his barnyard?
thanks!

Answer by checkley71(8403)

g+c=26 counting heads
4g+2c=72 counting feet
if we multiply the ehad count by -2 and add we get the following
-2g-2c=-52
4g+2c=72
--------------------------------
2g=20
g=10 count of goats thus
10+c=26
c=26-10
c=16 count of chickens
proof
4*10+2*16=72
40+32=72
72=72